OpenStudy (anonymous):

factor (x^3-1)/(x-1)

OpenStudy (anonymous):

You might find this a useful formula x³ + 1 = (x + 1)(x² - x + 1) x³ + y³ = (x + y) (x² - xy + y²)

OpenStudy (anonymous):

ill give you the anwer if you give me a medal lol:)

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

x^3=1 x=(1)^(1/3) x=1 x/x-1 you got it from here or you need me to answer it

OpenStudy (anonymous):

1/1-1

OpenStudy (anonymous):

Answer is 1